Change to standard equation for a circle.((x-x1)^2+(y-y1)^2=r^2) Find the center and radius. x^2-6x+9+y^2-4x+4=16
Answer by lwsshak3(11628)

Change to standard equation for a circle.((x-x1)^2+(y-y1)^2=r^2) Find the center and radius. x^2-6x+9+y^2-4x+4=16
***
Standard form of equation for a circle: (x-h)^2+(y-k)^2=r^2, (h,k)=(x,y) coordinates of center, r=radius.
..
x^2-6x+9+y^2-4x+4=16
x^2-10x+y^2=3
complete the square:
(x^2-10x+25)+y^2=3+25
(x-5)^2+y^2=28
center:(5,0)
radius: √28≈5.29